# MemeCLIP: Leveraging CLIP Representations for Multimodal Meme Classification

### v1

@@ 3.4 Statistics and Inter-Annotator Agreement
Table 2 shows the distribution of images in PrideMM across all class labels. For the hate detection task, the dataset has a balanced distribution of binary labels. The target detection task exhibits an imbalanced distribution. Given the context of this study, most hateful images convey undirected hate or are targeted toward communities, with low occurrences of hate against individuals and organizations. For the stance detection task, the number of images is well-balanced along three labels. On the other hand, as memes are often meant to be humorous, the majority of the images in the dataset are humorous. We provide a detailed analysis of PrideMM’s text content by using topic modeling in Appendix A.1.
We used the Fleiss’ Kappa (${\kappa}$) (Falotico and Quatto, 2015) as a statistical measure to assess the inter-annotator agreement across all four tasks. For Task A (Hate Speech detection), the Fleiss’(${\kappa}$) was 0.66/0.74 in the dry run and final phase respectively, for Task B (Target detection), ${\kappa}$ was 0.68/0.81, for Task C (Stance detection), ${\kappa}$ was 0.62/0.75, and for Task D (Humor detection), ${\kappa}$ was 0.60/0.74. The increase in the ${\kappa}$ score from the dry run phase to the final phase across all tasks reflects the effectiveness of the 3-phase annotation schema.
@@
Figure 2: An overview of our proposed framework, MemeCLIP. We use frozen image and text CLIP encoders to create representations for each image-text pair. These representations are passed through linear layers to disentangle the modalities in CLIP’s shared embedding space. We implement Feature Adapters with residual connections for each modality to prevent overfitting. We use a cosine classifier to make MemeCLIP more robust to imbalanced data. We initialize classifier weights by using Semantic-Aware Initialization to further improve performance.
@@ 4 Methodology
where $L^{proj}_{I}$ and $L^{proj}_{T}$ represent the image and text projection layers respectively.
Feature Adapters.  Since CLIP is pre-trained on an extensive amount of data, it may exhibit symptoms of overfitting when applied to smaller datasets for downstream tasks. Inspired by Gao et al. (2024), we adopt lightweight Feature Adapters for both image and text modalities to learn the features of the new data while retaining CLIP’s prior knowledge. We further utilize residual connections to integrate prior image and text projections to the outputs of the Adapters, allowing our model to balance the knowledge of the fine-tuned adapter and the disentangled image and text projections. We use a residual ratio $\alpha$ to maintain harmony between these two modules. With the image and text Feature Adapters $A_{I}$ and $A_{T}$ respectively, the final unimodal representations $F_{I},F_{T}\in\mathbb{R}^{1024}$ are obtained as follows:
@@
Classification.  For classification, we employ a cosine classifier Liu et al. (2020) that is more robust to biases in prediction under data imbalances. Following Shi et al. (2023), we adopt Semantic-Aware Initialization (SAI) to initialize the weights of this classifier by exploiting the semantic knowledge within the text encoder of CLIP. We encode class labels by using the prompt "A photo of {LABEL}" into $F_{class}\in\mathbb{R}^{n\times 1024}$ where n is the number of classes, and use $F_{class}$ to initialize the classifier weight $W_{class}$. During training, the predicted logit $Z$ for a class $x$ is calculated as follows:
@@
Table 3: Classification performance of models on the PrideMM dataset. The results are in the form of Mean ± Standard Deviation. Performance is reported across three evaluation metrics: Accuracy, AUROC (Macro), and F1-Score (Macro). The best performance is highlighted in bold.
@@ 5 Experimental Results
Table 3 and Table 4 show our experimental results for the PrideMM and HarMeme datasets respectively. We pre-define train/validation/test splits in the ratio 85/5/10 respectively for PrideMM, and use the pre-defined split for the HarMeme dataset. We conduct experiments on unimodal and multimodal baseline methods and previous frameworks proposed for multimodal hate classification. We conduct each experiment on three random seeds and report the Mean and Standard Deviation (±) values for Accuracy, AUC (Macro), and F1-Score (Macro). We use ViT-L/14 as the image encoder for all CLIP-based methods except for MOMENTA, which uses ViT-B/32 as the backbone for its CLIP model by default. For CLIP, we use concatenation to fuse unimodal feature representations. Further implementation details are outlined in A.3.
@@ 5.1 PrideMM Dataset
Multimodal Methods.  We tested the performance of the multimodal methods CLIP Radford et al. (2021), CLIP-Adapter Gao et al. (2024), MOMENTA Pramanick et al. (2021b), HateCLIPper Kumar and Nandakumar (2022), ISSUES Burbi et al. (2023), and our framework, MemeCLIP. MemeCLIP outperforms the baseline CLIP model and previously proposed multimodal methods across all metrics in the hate and humor classification tasks. Our model performs particularly well in the four-class target classification task which has merely half the number of samples as the other tasks and harbors a heavy class imbalance, demonstrating the model’s robustness to overfitting and data imbalance settings. Smaller frameworks with a lower number of parameters such as the baseline CLIP, CLIP-Adapter, HateCLIPper, and MemeCLIP perform optimally in this task with CLIP-Adapter showing the highest AUC, while the remaining methods show symptoms of overfitting. In the stance classification task, HateCLIPper surpasses MemeCLIP in accuracy, while the latter shows a higher AUC and F1-Score. While the CLIP-Adapter model is similar to MemeCLIP, it uses a single feature adapter as individual feature adapters for both modalities may carry redundant information from the unimodal encoders to the classifier; however, MemeCLIP outperforms this model in most tasks by using feature adapters for both image and text modalities, which may signify that separate layers for each modality are more effective at encoding memes that may contain visual and language content that convey different meanings individually and when put together.
@@
Table 4: Classification performance of models on the HarMeme dataset. The results are in the form of Mean ± Standard Deviation. Performance is reported across three evaluation metrics: Accuracy, AUROC (Macro), and F1-Score (Macro). The best performance is highlighted in bold.
@@ 5.2 HarMeme Dataset
To test the generalizability of MemeCLIP across other meme datasets, we perform all the experiments on the HarMeme dataset Pramanick et al. (2021a), which consists of real-world hateful memes shared on social media in the context of COVID-19. Similar to previous studies, we find that the performance of the multimodal models surpasses the unimodal models’ performance by a decent margin. MemeCLIP outperforms the other multimodal frameworks in this dataset, which demonstrates its effectiveness over the state-of-the-art baselines.
@@ 5.3 Ablation Study
We conduct a systematic ablation study to assess the contribution of each component of MemeCLIP to its performance. The results of our ablation experiments are presented in Table 5. We start with the CLIP ViT-L/14 model and gradually integrate each external module of MemeCLIP. The rise in performance when the projection layers are applied highlights the importance of adapting CLIP’s embedding spaces to our downstream task by disentangling the image and text representations. While the introduction of Feature Adapters initially leads to a temporary dip in F1-Score, it ultimately enables our model to produce more refined image and text representations due to the added learnable parameters. Replacing the linear classifier with a cosine classifier boosts performance by modulating weight updates with a static scaling factor. Finally, Semantic-aware initialization completes MemeCLIP by initializing classifier weights according to the semantic differences in class labels encoded by CLIP’s text encoder, enhancing generalization further.
Table 5: Ablation experiments performed on MemeCLIP. The results are in the form of Mean ± Standard Deviation. PL, FA, CC, and SAI denote Projection Layers, Feature Adapters, Cosine Classifier, and Semantic-Aware Initialization respectively. The last line represents the complete framework. The best results are highlighted in bold.
@@ 5.4 Comparison with GPT-4
Table 6: Performance comparison between MemeCLIP and the GPT-4 model provided by Microsoft Copilot on hate classification for PrideMM and harm classification for HarMeme.
@@ 6 Conclusion
In this work, we release PrideMM, a multimodal and multi-aspect dataset comprising 5,063 memes related to the LGBTQ+ movement, addressing a serious gap in data resources in this domain. This dataset provides memes annotated across three aspects and one sub-class, allowing for greater flexibility in the establishment of ethical guidelines by social media policymakers. We further introduce MemeCLIP, a lightweight yet effective framework to harness CLIP’s knowledge for multimodal meme classification while not relying on extraneous models to create augmented data for training. With each module of MemeCLIP, we tried to address pervasive problems in the domain of hateful image classification. Specifically, we used a cosine classifier to counter class imbalances, which are common in multi-aspect datasets with many classes for each aspect. We also utilize feature adapters to mitigate overfitting as datasets in this domain have a relatively small scale. We further use linear projection layers to dissociate the image and text modalities in the representations created by the pre-trained CLIP’s encoders as the pre-training dataset of CLIP majorly consists of image-text pairs that convey the same overall meaning, which may not be the case with creative or sarcastic memes. We show that CLIP, one of the most basic multimodal models, when combined with lightweight additional modules, can compete with or outperform models that require extraneous models to create augmented data, or even VLMs such as GPT-4. Our work is grounded on the importance of a data-centric approach to solving problems in this domain, rather than creating larger frameworks and incorporating extraneous data. Our dataset endeavors to foster a deeper understanding of online interaction, community-building, and social change, whereas our framework is a step toward effective content moderation that helps create an inclusive, diverse, and equitable digital environment for all.
@@
Biases.  Due to the size of the dataset and the number of annotation tasks, we acknowledge that some data samples may be misclassified, and not all annotators would agree on the same labels for a given sample. Further, due to the subjective nature of this topic, unintentional biases in data distribution and annotation may exist. We endeavored to minimize such occurrences by using comprehensive annotation guidelines and a three-phase annotation schema.
@@
Reproducibility Statement.  We provide implementation details and hyperparameter configurations for all the implemented models in Appendix A.3. The PrideMM dataset, source code for MemeCLIP, and MemeCLIP pre-trained weights are publicly available at https://github.com/SiddhantBikram/MemeCLIP.
Environmental Impact.  Leveraging hardware such as GPUs to train deep learning models is known to have a significant environmental footprint, primarily due to their high energy consumption resulting in carbon emissions. To mitigate the environmental impact of our research, we adopted a fine-tuning technique for pre-trained models. This method allowed our models to generalize faster on our dataset resulting in fewer computational resources consumed. Once these models are trained using GPUs, they can be loaded on relatively lightweight CPUs for inference purposes, attenuating potential environmental consequences.
@@
The proposed dataset PrideMM encompasses memes posted from 2020 to 2024, representing a snapshot of social media interactions only within this specific period, which may fail to capture the dynamics of the LGBTQ+ movement on social media over an extended timeframe. Further, due to the subjective nature of the LGBTQ+ Movement, the size of the dataset, and the number of annotators, the annotation process is inherently prone to biases. While we include images from multiple sources in our dataset, we acknowledge the limited scope of our dataset to these platforms. The intricate nature of memes may not be completely captured by our aspects and more domain-specific aspects may be used to capture the context better. Due to the limited size and availability of labeled datasets in this domain, supervised frameworks such as MemeCLIP may be surpassed by unsupervised methods such as LLMs for content moderation as the capabilities and model size of LLMs progress. Further, the model may exhibit biased predictions due to unintentionally biased data in such datasets.
@@ A.1 Topic Modeling
We applied the Sparse Additive Generative Models of Text (SAGE) Eisenstein et al. (2011) topic modeling technique to identify noteworthy words across various class labels within our dataset. We set the hyperparameters max_vocab_size to 1000 and base_rate_smoothing to 1. Tables 8 and 7 present the most notable words for each class as identified by SAGE along with their corresponding salience scores. Among hate targets, words such as "rowling", "shapiro", "conservative", "gays", "bethesda", and "corporations" are assigned high scores by SAGE, helping identify the most targeted entities for each label. Within samples labeled Support, words such as "comfortable", "expression", and "supportive" hold relevance as they convey acceptance and support.
@@
Hate Target
No Hate Hate Undirected Individual Community Organization
@@ A.2 Data Filtering
  • Low-Quality Images: We discard images that were highly distorted, blurred, or degraded. We also removed images with illegible text.
@@ A.3 Implementation Details
We conduct all our experiments on Pytorch 2.1.2 and an NVIDIA Tesla T4 GPU with 16 GB of dedicated memory. We set the batch size to 16 and trained each model for 10 epochs while monitoring validation AUROC to save the best model for each run. Under these settings, training and validating MemeCLIP for one epoch takes 12 minutes and occupies 7 GB of dedicated memory.
We empirically find the most optimal learning rate for each model. We use a learning rate of $10^{-5}$ for ViT-L/14 and CLIP Image-Only. We use a learning rate of $5\times 10^{-5}$ for BERT and CLIP Text-Only. We use a learning rate of $10^{-3}$ for CLIP and CLIP-Adapter. For MOMENTA, HateCLIPper, and ISSUES, we use the default settings set by the authors. For MemeCLIP, we set the learning rate to $10^{-4}$. We set the scaling factor $\sigma$ for the cosine classifier to 30, and the residual ratio $\alpha$ for the Feature Adapters to 0.2.
We implemented CLIP, CLIP Image-Only, and CLIP Text-Only models by using OpenAI’s official CLIP library [^5] https://github.com/openai/CLIP. Our framework, MemeCLIP was built upon the base CLIP model provided by this library. We implemented the Visual Transformer model by using the timm library provided by Huggingface [^6] https://github.com/huggingface/pytorch-image-models. We implemented BERT by using the Huggingface Transformers library [^7] https://github.com/huggingface/transformers. We obtained the code released by the authors of CLIP-Adapter [^8] https://github.com/gaopengcuhk/CLIP-Adapter, MOMENTA [^9] https://github.com/LCS2-IIITD/MOMENTA, HateCLIPper [^10] https://github.com/gokulkarthik/hateclipper, and ISSUES [^11] https://github.com/miccunifi/ISSUES to test their respective methods. The total number of parameters for each method is listed in Table 9.

### v2

@@ 3.4 Statistics and Inter-Annotator Agreement
Table 2 shows the distribution of images in PrideMM across all class labels. For the hate detection task, the dataset has a balanced distribution of binary labels. The target classification task exhibits a heavily imbalanced distribution. Given the context of this study, most hateful images convey undirected hate or are targeted toward communities, with a low frequency of hate against individuals and organizations. For the stance classification task, the number of images is well-balanced across three labels. On the other hand, as memes are often meant to be humorous, the majority of the images in the dataset are annotated to humor. We use topic modeling to analyze PrideMM’s text content, and the results are presented in Appendix A.1.
We used the Fleiss’ Kappa (${\kappa}$) (Falotico and Quatto, 2015) as a statistical measure to assess the inter-annotator agreement across all four tasks. For Task A (Hate Speech detection), ${\kappa}$ was 0.66/0.74 in the dry run and final phase respectively, for Task B (Target detection), ${\kappa}$ was 0.68/0.81, for Task C (Stance detection), ${\kappa}$ was 0.62/0.75, and for Task D (Humor detection), ${\kappa}$ was 0.60/0.74. The increase in ${\kappa}$ from the dry run phase to the final phase across all tasks reflects the effectiveness of the 3-phase annotation schema.
@@
Figure 2: An overview of our proposed framework, MemeCLIP. We use frozen CLIP image and text encoders to create representations for each image-text pair. These representations are passed through linear layers to disentangle the modalities in CLIP’s shared embedding space. We implement Feature Adapters with residual connections for each modality to prevent overfitting. We use a cosine classifier to make MemeCLIP more robust to imbalanced data. We initialize classifier weights by using Semantic-Aware Initialization to further improve performance.
@@ 4 Methodology
Here, $L^{proj}_{I}$ and $L^{proj}_{T}$ represent the image and text projection layers respectively.
Feature Adapters.  Since CLIP is pre-trained on an extensive amount of data, it may exhibit symptoms of overfitting when applied to smaller datasets for downstream tasks. Inspired by Gao et al. (2024), we adopt lightweight Feature Adapters for both image and text modalities to learn the features of new data while retaining CLIP’s prior knowledge. We further utilize residual connections to integrate prior image and text projections with the outputs of the Adapters, allowing our model to balance the knowledge of the fine-tuned adapter and the disentangled image and text projections. We use a residual ratio $\alpha$ to maintain harmony between these two modules. With the image and text Feature Adapters $A_{I}$ and $A_{T}$ respectively, the final unimodal representations $F_{I},F_{T}\in\mathbb{R}^{1024}$ are obtained as follows:
@@
Classification.  For classification, we employ a cosine classifier Liu et al. (2020) that is robust to biases in prediction under class imbalances. Following Shi et al. (2023), we adopt Semantic-Aware Initialization (SAI) to initialize the weights of this classifier by exploiting the semantic knowledge held within the text encoder of CLIP. We encode class labels by using the prompt "A photo of {LABEL}" into $F_{class}\in\mathbb{R}^{n\times 1024}$ where n is the number of classes. We use $F_{class}$ to initialize the classifier weight $W_{class}$. During training, the predicted logit $Z$ for a class $x$ is calculated as follows:
@@
Table 3: Classification performance of methods on the PrideMM dataset. The results are in the form of Mean ± Standard Deviation. Performance is reported across three evaluation metrics: Accuracy, AUROC (Macro), and F1-Score (Macro). The best performance is highlighted in bold.
@@ 5 Experimental Results
Table 3 and Table 4 show our experimental results for the PrideMM and HarMeme datasets respectively. We pre-define train/validation/test splits in the ratio 85/5/10 respectively for PrideMM, and use the pre-defined split for the HarMeme dataset. We conduct experiments on unimodal and multimodal baseline methods, and previous frameworks proposed for multimodal meme classification. We conduct each experiment on three random seeds and report the Mean and Standard Deviation (±) values for Accuracy, AUC (Macro), and F1-Score (Macro). We use ViT-L/14 as the image encoder for all CLIP-based methods except for MOMENTA, which uses ViT-B/32 as the backbone for its CLIP model by default. For CLIP, we use concatenation to fuse the unimodal feature representations. Further implementation details are outlined in A.3.
@@ 5.1 PrideMM Dataset
Multimodal Methods.  We tested the performance of the multimodal methods CLIP Radford et al. (2021), CLIP-Adapter Gao et al. (2024), MOMENTA Pramanick et al. (2021b), HateCLIPper Kumar and Nandakumar (2022), ISSUES Burbi et al. (2023), and our framework, MemeCLIP. MemeCLIP outperforms the baseline CLIP model and previously proposed multimodal methods across all metrics in the hate and humor classification tasks. Our model performs particularly well in the four-class target classification task, which has less than half the number of samples as the other tasks and harbors a heavy class imbalance, demonstrating the model’s robustness to overfitting on majority classes in imbalanced datasets. Smaller frameworks with a lower number of parameters such as the baseline CLIP, CLIP-Adapter, HateCLIPper, and MemeCLIP perform optimally in this task with CLIP-Adapter showing the highest AUC, while the remaining methods show symptoms of overfitting. In the stance classification task, HateCLIPper surpasses MemeCLIP in accuracy, while the latter shows a higher AUC and F1-Score. While the CLIP-Adapter model is similar to MemeCLIP, it uses a single feature adapter as individual feature adapters for both modalities may carry redundant information from the unimodal encoders to the classifier; however, MemeCLIP outperforms this model in most tasks by using feature adapters for both image and text modalities, which may signify that separate layers for each modality are more effective at encoding memes that may contain visual and language content that convey different meanings individually and combined.
@@
Table 4: Classification performance of methods on the HarMeme dataset. The results are in the form of Mean ± Standard Deviation. Performance is reported across three evaluation metrics: Accuracy, AUROC (Macro), and F1-Score (Macro). The best performance is highlighted in bold.
@@ 5.2 HarMeme Dataset
To test the generalizability of MemeCLIP across other meme datasets, we perform experiments on the HarMeme dataset Pramanick et al. (2021a), which consists of real-world hateful memes shared on social media in the context of COVID-19. Similar to previous studies, we find that the performance of the multimodal models surpasses the unimodal models’ performance by a decent margin. MemeCLIP outperforms the other multimodal frameworks in this dataset, demonstrating its effectiveness over the state-of-the-art baselines.
@@ 5.3 Ablation Study
We conduct a systematic ablation study to assess the contribution of each component of MemeCLIP towards its performance. The results of our ablation experiments are presented in Table 5. We start with the CLIP ViT-L/14 model and gradually integrate each external module of MemeCLIP. The rise in performance when the projection layers are applied signifies the importance of adapting CLIP’s embedding spaces to our downstream task by disentangling the image and text representations. While the introduction of Feature Adapters initially leads to a temporary dip in F1-Score, it ultimately enables our model to produce more refined image and text representations due to the added learnable parameters. Replacing the linear classifier with a cosine classifier boosts performance by modulating weight updates with a static scaling factor. Finally, Semantic-aware initialization completes MemeCLIP by initializing classifier weights according to the semantic differences in class labels encoded by CLIP’s text encoder, enhancing generalization further.
Table 5: Ablation experiments performed on MemeCLIP using the hate detection task of the PrideMM dataset. The results are in the form of Mean ± Standard Deviation. PL, FA, CC, and SAI denote Projection Layers, Feature Adapters, Cosine Classifier, and Semantic-Aware Initialization respectively. The last line represents the complete framework. The best results are highlighted in bold.
@@ 5.4 Comparison with GPT-4
Table 6: Performance comparison between MemeCLIP and the GPT-4 model provided by Microsoft Copilot on hate classification for PrideMM and harm classification for HarMeme. The best results are highlighted in bold.
@@ 6 Conclusion
In this work, we release PrideMM, a multimodal and multi-aspect dataset comprising 5,063 memes related to the LGBTQ+ movement, addressing a serious gap in data resources in this domain. This dataset provides memes annotated across three aspects and one sub-class, allowing for greater flexibility in the establishment of ethical guidelines by social media policymakers. We further introduce MemeCLIP, a lightweight yet effective framework to harness CLIP’s knowledge for multimodal meme classification while not being reliant on extraneous models to create augmented data for training.
With each module of MemeCLIP, we tried to address pervasive problems in the domain of hateful image classification. Specifically, we used a cosine classifier to counter class imbalances, which are common in multi-aspect datasets with many classes for each aspect. We utilize feature adapters to mitigate overfitting as datasets in this domain have a relatively small scale. We further use linear projection layers to dissociate the image and text modalities in the representations created by the pre-trained CLIP encoders as the pre-training dataset of CLIP majorly consists of image-text pairs that convey the same overall meaning, which may not be the case with creative or sarcastic memes. We show that CLIP, one of the most basic multimodal models, when combined with lightweight additional modules, can compete with or outperform models that require extraneous models to create augmented data, or even VLMs such as GPT-4. Our work is grounded on the importance of a data-centric approach to solving problems in this domain, rather than creating larger frameworks and incorporating extraneous data.
Our dataset endeavors to foster a deeper understanding of online interaction, community-building, and social change, whereas our framework is a step toward effective content moderation that helps create an inclusive, diverse, and equitable digital environment for all.
@@
Biases.  Due to the size of the dataset and the number of annotation tasks, we acknowledge that some data samples may be misclassified, and not all annotators would agree on the same labels for a given sample. Further, due to the subjective nature of this topic, unintentional biases in the dataset’s distribution and annotation may exist. We endeavored to minimize such occurrences by using comprehensive annotation guidelines and a three-phase annotation schema.
@@
Reproducibility Statement.  We provide implementation details and hyperparameter configurations for all the implemented models in Appendix A.3. The PrideMM dataset, source code for MemeCLIP, and MemeCLIP’s pre-trained weights are publicly available at https://github.com/SiddhantBikram/MemeCLIP.
Environmental Impact.  Leveraging hardware such as GPUs to train deep learning models is known to have a significant environmental footprint, primarily due to their high energy consumption resulting in carbon emissions. To mitigate the environmental impact of our research, we adopted a fine-tuning technique for pre-trained deep-learning models. This method allowed our models to generalize faster on new datasets, resulting in fewer computational resources consumed. Once these models are trained using GPUs, they can be loaded on relatively lightweight CPUs for inference purposes, attenuating potential environmental consequences.
@@
The proposed dataset PrideMM encompasses memes posted from 2020 to 2024, representing a snapshot of social media interactions within this specific period, which may fail to capture the dynamics of the LGBTQ+ movement on social media over an extended timeframe. Further, due to the subjective nature of the LGBTQ+ Movement, the size of the dataset, and the number of annotators, the annotation process is inherently prone to biases. While we include images from multiple sources in our dataset, we acknowledge the limited scope of our dataset compared to the vast number of social media platforms. Further, the intricate nature of memes may not be completely captured by our aspects, and more domain-specific aspects may be used to capture the context better. Due to the limited size and availability of labeled datasets in this domain, supervised frameworks such as MemeCLIP may be surpassed by unsupervised methods such as LLMs for content moderation as the capabilities and model size of LLMs progress. Finally, MemeCLIP and other models trained on PrideMM and similar datasets may exhibit biased predictions due to the presence of unintentionally biased data.
@@ A.1 Topic Modeling
We applied the Sparse Additive Generative Models of Text (SAGE) Eisenstein et al. (2011) topic modeling technique to identify noteworthy words across various class labels within our dataset. We set the hyperparameters max_vocab_size to 1000 and base_rate_smoothing to 1. Tables 7 and 8 present the most notable words for each class as identified by SAGE along with their corresponding salience scores. Among hate targets, words like ‘rowling’, ‘shapiro’, ‘conservative’, ‘gays’, ‘bethesda’, and ‘corporations’ are assigned high scores by SAGE, helping identify the most targeted entities for each label. Within samples labeled Support, words such as ‘comfortable’, ‘expression’, and ‘supportive’ hold relevance as they convey acceptance and support.
Hate Target
No Hate Hate Undirected Individual Community Organization
@@ A.2 Data Filtering
  • Low-Quality Images: We discarded images that were highly distorted, blurred, or degraded. We also removed images with illegible text.
@@ A.3 Implementation Details
We conducted all our experiments on Pytorch 2.1.2 combined with an NVIDIA Tesla T4 GPU with 16 GB of dedicated memory. We set the batch size to 16 and trained each model for 10 epochs while monitoring validation AUROC to save the best model for each run. Under these settings, training and validating MemeCLIP for one epoch takes 12 minutes and occupies 7 GB of dedicated memory.
We empirically found the most optimal learning rate for each model. We used a learning rate of $10^{-5}$ for ViT-L/14 and CLIP Image-Only. We used a learning rate of $5\times 10^{-5}$ for BERT and CLIP Text-Only. We used a learning rate of $10^{-3}$ for CLIP and CLIP-Adapter. For MOMENTA, HateCLIPper, and ISSUES, we used the default settings set by their respective authors. For MemeCLIP, we set the learning rate to $10^{-4}$. We set the scaling factor $\sigma$ for the cosine classifier to 30, and the residual ratio $\alpha$ for the Feature Adapters to 0.2.
Our framework, MemeCLIP, was built upon the base CLIP model provided by OpenAI’s official CLIP library [^5] https://github.com/openai/CLIP. We also implemented CLIP, CLIP Image-Only, and CLIP Text-Only models by using this library. We implemented the Visual Transformer model by using the timm library provided by Huggingface [^6] https://github.com/huggingface/pytorch-image-models. We implemented BERT by using the Huggingface Transformers library [^7] https://github.com/huggingface/transformers. We obtained the code released by the authors of CLIP-Adapter [^8] https://github.com/gaopengcuhk/CLIP-Adapter, MOMENTA [^9] https://github.com/LCS2-IIITD/MOMENTA, HateCLIPper [^10] https://github.com/gokulkarthik/hateclipper, and ISSUES [^11] https://github.com/miccunifi/ISSUES to test their respective methods. The total number of parameters for each method is listed in Table 9.